OpenStudy (anonymous):

Solve the equation and check the solution: \[\dfrac{a}{(a-6) (a+6)} + \dfrac{2}{(a-6)} = \dfrac{1}{(a+6)}\]

OpenStudy (whpalmer4):

Two choices: put all the fractions over a common denominator, or clear the fractions by multiplying through by \((a-6)(a+6)\)

OpenStudy (anonymous):

\[\dfrac{a}{a^2-36}+\dfrac{2}{a-6}= \dfrac{1}{a+6}\] \[\dfrac{a}{(a-6) (a+6)} + \dfrac{2}{(a-6)} = \dfrac{1}{(a+6)}\] \[\dfrac{a}{(a-6)(a+6)} + \dfrac{2(a+6)}{(a-6)(a+6)} = \dfrac{(a-6)}{(a+6)(a-6)}\] \[\dfrac{a + 2(a+6) - (a-6)}{(a+6)(a-6)} = 0\] \[\dfrac{2+2a+16-a+6}{(a+6)(a-6)}=0\] \[\dfrac{a+24}{(a+6)(a-6)}= 0\] \[a+24 = 0\] \[a = -24 \] My textbook says the answer should be -9 what did I do wrong?

OpenStudy (anonymous):

In 4th last step, you have done a silly mistake my dear.The correct equation will be : (a+2a+12-a+6)/(a-6)(a+6)=0 (2a+18)=0 2a=-18 a=-9 which is the answer suggested by your textbook

OpenStudy (whpalmer4):

I would clear the denominators, giving me\[a + 2(a+6) = a-6\]\[3a + 12 = a-6\]\[2a=-18\]\[a=-9\]

OpenStudy (whpalmer4):

Your mistake was taking \(2(a+6)\) and making \(2a+16\) out of it instead of \(2a+12\)